OpenStudy (anonymous):

what are two equivalent for 4/11

OpenStudy (anonymous):

8/22,12/33,16/44,20/55,24/66,28/77,32/88,36/99,40/110 those are fractions that are equivalent to 4/11

OpenStudy (anonymous):

o 4/11 X 3/3 = 12/33 or 4/11 X 10/10 = 40/110